Ops/Comm Scheduling meeting

Location : CR21 in R1 @ 15:30 pm on August 16th

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

  • AOB
  • DONM : 23 August in CR21

Henry, Pierrick, Paul K, Colin W, Paul H, Victoria B, Vicky B, Josef B, John G, Ken L

DC report :

Continuing water system. FC and SSD refilled last Wednesday and that went well with high helium fill efficiency of 70% or so.
SSD condensed to 98%. Used 250 L to fill 180 L. John has brought a 5th AC unit into the Hall and a smaller chiller to cool the
transfer lines.

There's an issue with the AC units in RR2. They were in an alarm state on Monday. They turned one on again and the temp is still
quite high. John will check this.

Water system:

Primary system up in the roof. ICS will come this week and do the controls. If they do this, we can change on Monday/Tuesday.
Daresbury people come back on Tuesday for more powering. Daresbury need to be able to run the conventional magnets next week.
Waterworks on the magnet circuit down in the trench. A conductivity meter has been acquired but there's still the problem of
scheduling ISIS effort in implementing the temperature control.

The lads are starting to work on the AC tomorrow. Will concentrate on #1 and #4 and will install #5. THey will be in there the
rest of the week. Linde fridge can be brought up with #4 is ready so by end of the week. Vicky will plan to start next Monday
on cooling down the Linde fridge.

Josef: we need instructions on how to swap between chiller and roof water.
Colin : we should tell shifters what valves to close if there is a problem with the water.


Hall open for the AC work until Tuesday/Wednesday next week so no magnet work since Hall is open.
Daresbury are trying different configurations today of the resistor-diode path. It seems the fluctuations we are seeing in the
voltage come from the diodes. Suspicion is that the fluctuations come from the power supply.

Hydrogen system service:

What do service the hydrogen system sensors. The company asked to come on the 15th September. 5-9/12 or 28th.

DL : Decay solenoid : Current calirbation first full week in September

MOdify item for SS contact:

Updated by Boyd, Steven over 7 years ago ยท 3 revisions